[Hallc_running] [New Logentry] RC daily update (8/24)
billlee at jlab.org
billlee at jlab.org
Mon Aug 24 14:35:12 EDT 2020
I am very happy to tell you that the target installation (Briana) was successful, the position alignments at oven cold and hot were acceptable (https://logbooks.jlab.org/entry/3833229). There is no leak observed in the reference cell (https://logbooks.jlab.org/entry/3833210). Before ramping to all magnets to full field (HMS to -4.5 GeV and SHMS to -7.5 GeV, note that HB is already ramped), the initial NMR result at 2pm is 8%. (https://logbooks.jlab.org/entry/3833315).
In terms of other Hall activities,
1. Simona did a lot great work on the Cherenkov test, also, she was able to certify SHMS's moving ability between 11 to 18 degrees this morning.
2. The main air compressor for the target failed over the weekend, we are on the backup compressor right now. Walter arranged to rent a diesel air compresses ready to be used for the experiment as a spare.
3. Ethan and William pushed some changes in code to the target ladder control, these changes will take effect after rebooting the IOC. Hopefully, this will help with the target ladder vibration issue.
Planning ahead, we are looking at closing the target up before 6pm this afternoon, and lockup for beam soon after that. We are likely to take some reference cell/production data the over night. Moller measurement is tentatively planned for tomorrow.
>From the accelerator side, there is an issue with 1L13-7 cavity, which was zeroed this morning 11 am. PD is attempting to coordinate to have it fixed tomorrow (2 hour downtime projected). We have to decide if we would like to perform the moller measurement after or wait after the 1L13-7 repair to start. Note that the not fixing the issue increases the RF related trip count by 1.5 /hour.
We will have a meeting at 4:15pm today. To further discussion on the running plan.
This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/3833344
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Hallc_running